### v1

# TwinVLA: Data-Efficient Bimanual Manipulation with Twin Single-Arm Vision-Language-Action Models

Abstract
Vision-language-action models (VLAs) trained on large-scale robotic datasets have demonstrated strong performance on manipulation tasks, including bimanual tasks. However, because most public datasets focus on single-arm demonstrations, adapting VLAs for bimanual tasks typically requires substantial additional bimanual data and fine-tuning. To address this challenge, we introduce TwinVLA, a modular framework that composes two copies of a pretrained single-arm VLA into a coordinated bimanual VLA. Unlike monolithic cross-embodiment models trained on mixtures of single-arm and bimanual data, TwinVLA improves both data efficiency and performance by composing pretrained single-arm policies. Across diverse bimanual tasks in real-world and simulation settings, TwinVLA outperforms a comparably-sized monolithic RDT-1B model without requiring any bimanual pretraining. Furthermore, it narrows the gap to state-of-the-art model, $\pi_{0}$ which rely on extensive proprietary bimanual data and compute cost. These results establish our modular composition approach as a data-efficient and scalable path toward high-performance bimanual manipulation, leveraging public single-arm data.
1 Introduction
Thanks to publicly available large-scale robotic datasets, vision-language-action models (VLAs) have shown impressive performance in single-arm robotic manipulation, generalizing across diverse tasks, objects, and environments (rt22023arxiv; oxe2024; kim2024openvla; black2024pi_0). However, extending these successes to bimanual manipulation remains challenging, as public bimanual datasets are scarce, and existing approaches often rely on large, proprietary datasets that require thousands of hours of data collection and curation (black2024pi_0), limiting reproducibility and progress.
Can we build strong bimanual VLAs without collecting or fine-tuning on large bimanual datasets by leveraging existing single-arm data? Recent cross-embodiment learning work typically trains monolithic models on multi-robot datasets (oxe2024) by employing embodiment-specific action decoders (octo_2023; doshi2024scalingcrossembodiedlearningpolicy; nvidia2025gr00tn1openfoundation) or shared, zero-padded action spaces (liu2024rdt; black2024pi_0). Although promising, differences in observation and action spaces introduce heterogeneity, forcing a single model to handle disparate action spaces, and monolithic training underutilizes the modular structure inherent to bimanual tasks.
A modular perspective on bimanual manipulation is supported by neuroscience: human bimanual manipulation is the coordination of arm-specific motor primitives rather than a single monolithic controller. Dedicated neural circuits, such as the Supplementary Motor Area (SMA) and the corpus callosum, orchestrate and synchronize the two arms (Sadato9667; Swinnen2002). Similar principles have proven effective in vision-language modeling, where interaction between modality-specific backbones improves its efficiency and effectiveness (liang2024mixtureoftransformerssparsescalablearchitecture).
Figure 1: Overview of TwinVLA. Inspired by humans’ two-arm coordination for bimanual manipulation, TwinVLA duplicates a VLM backbone pretrained on cross-embodiment single-arm data (Left) to form two arm-specific branches linked via Joint Attention (Right). Shared inputs (ego-centric views, language instructions) are routed via a mixture-of-experts (MoE) to improve computational efficiency. Only the VLM backbone is duplicated, keeping the increase in model size minimal.
Inspired by these insights, we propose TwinVLA, a modular architecture that operationalizes this coordination-centric view. Instead of training from scratch, TwinVLA leverages a pretrained single-arm VLA. Specifically, we first design a lightweight, compact single-arm VLA, which we call SingleVLA (Appendix˜A). We pretrain SingleVLA for single-arm manipulation on the OXE dataset (oxe2024). We then duplicate this SingleVLA and integrate the two “twin” instances through a lightweight coordination method. This design is highly data-efficient: it eliminates the need for a bimanual pretraining dataset and achieves strong performance with only a small amount of bimanual demonstrations for fine-tuning.
To integrate two SingleVLAs into a bimanual policy, TwinVLA utilizes a joint attention (liang2024mixtureoftransformerssparsescalablearchitecture) across the twin models, as illustrated in Figure˜1. This allows the twin SingleVLAs to exchange information and coordinate their actions, while preserving their pretrained capabilities. This approach is made feasible without significant overhead, as we duplicate only the VLM backbone and utilize a Mixture-of-Experts (MoE) to efficiently manage shared inputs. In contrast to monolithic cross-embodiment models (liu2024rdt; octo_2023; doshi2024scalingcrossembodiedlearningpolicy), our approach yields better performance and data efficiency, significantly reducing the need for large-scale bimanual data collection and compute.
We evaluate TwinVLA across a broad range of environments, including a complex, long-horizon real-world task and a diverse suite of bimanual manipulation tasks in simulations. Despite leveraging only public single-arm data and limited bimanual fine-tuning data, TwinVLA achieves performance comparable to state-of-the-art bimanual policies.
In summary, our main contributions are threefold:
  • We propose a novel modular architecture for bimanual manipulation that integrates two copies of a pretrained SingleVLA with a lightweight coordination method based on joint attention with MoE, enabling synchronized two-arm control.
  • We present a data-efficient paradigm that adapts our twin architecture into a capable bimanual policy for a target task by fine-tuning on only a small bimanual dataset, crucially without requiring additional pretraining, thereby eliminating the need for large-scale bimanual data.
  • Through extensive experiments across real and simulated bimanual tasks, TwinVLA matches or surpasses state-of-the-art models trained on far larger bimanual data and compute.
Together, these findings identify our modular SingleVLA composition approach as a scalable, efficient path to high-performance bimanual manipulation.
(a) Data efficiency comparison
(b) Compute efficiency comparison
Figure 2: (a) Data efficiency. While RDT-1B and $\pi_{0}$ use 1M + single-arm data with sizable bimanual data, TwinVLA uses $\sim$0.5M single-arm data and only 50 target bimanual data. (b) Compute efficiency. RDT-1B and $\pi_{0}$ require high compute (exceeding 1,000 H100 GPU-days), whereas TwinVLA achieves higher or comparable performance at substantially lower compute (25 H100 GPU-days).
2 Related Work
Vision-Language-Action models (VLAs) represent a promising approach to robotic control. They adapt large pretrained Vision-Language Models (VLMs) (liu2023visualinstructiontuning; karamcheti2024prismaticvlmsinvestigatingdesign; chen2024internvlscalingvisionfoundation) to robotic control (oxe2024; khazatsky2024droid) by fine-tuning on action-labeled data. This process enables them to translate language instructions and visual context into grounded actions. Early VLA models explore various strategies for connecting VLMs with action generation (ahn2022icanisay). RT-2 (rt22023arxiv) tokenizes actions as part of the model’s output vocabulary, whereas RoboFlamingo (li2024visionlanguagefoundationmodelseffective) uses a separate continuous action head. Recent work, such as $\pi_{0}$ (black2024pi_0) and CogAct (li2024cogactfoundationalvisionlanguageactionmodel), further integrates a denoising policy with a VLM to improve action execution accuracy.
Bimanual manipulation policies are essential for enabling robots to perform complex tasks that require coordinated two-handed control, such as folding laundry (bersch2011bimanual; avigal2022speedfolding), assembling parts (stavridis2018bimanual), or wiping the plate black2025pi; chi2024universalmanipulationinterfaceinthewild. Learning effective bimanual policies is challenging due to high-dimensional, tightly coupled action spaces and the scarcity of high-quality bimanual demonstrations (lee2020learning; xie2020deepimitationlearningbimanual). Consequently, specialist methods, such as Diffusion policy (chi2023diffusion) and ACT (zhao2023learning), trained only on target-task demonstrations, struggle on precise, long-horizon bimanual tasks (black2024pi_0).
To mitigate data scarcity, prior work has largely pursued a straightforward approach: training a single, unified model through large-scale bimanual data collection and computationally intensive pretraining. RDT-1B (liu2024rdt) is fine-tuned on over 66K bimanual episodes after extensive pretraining on the large-scale OXE dataset (oxe2024), reportedly requiring a month on $48$ H100 GPUs. Similarly, $\pi_{0}$ (black2024pi_0) relies on a $10,000$-hour proprietary dataset with substantial computational cost. Moreover, because their data are not publicly accessible, reproducibility and broader adoption are limited.
In contrast to such monolithic cross-embodiment policies and compute-heavy bimanual pretraining, our approach adopts a modular, coordination-centric design. We first train a SingleVLA on large-scale public single-arm data, duplicate the pretrained SingleVLA and couple them via joint attention, and then fine-tune it on bimanual tasks—allowing each stage to benefit from the most suitable data (see Figure˜2). This composition-based approach avoids bimanual pretraining, requires only a small amount of bimanual fine-tuning, better preserving the capabilities of single-arm policies, and improves data and compute efficiency.
3 Preliminaries
This paper aims to develop a data-efficient framework for learning bimanual manipulation policies by building upon pretrained single-arm Vision-Language-Action (SingleVLA) models. This section formalizes the single-arm and bimanual settings, and describes the conditional flow-matching objective for training action heads of VLAs.
3.1 Formulating the Bimanual VLA Policy
Our goal is to extend a pretrained SingleVLA $\pi_{\text{single}}$ into a bimanual policy $\pi_{\text{twin}}$ applicable to target bimanual tasks. A VLA $\pi(A_{t}\mid o_{t})$ predicts an action chunk $A_{t}=(a_{t},a_{t+1},\dots,a_{t+T-1})$ of length $T$ from an observation $o_{t}$. For single-arm manipulation, the observation $o^{\text{{single}}}_{t}=\left((l,I_{\text{ego}})_{t},(I_{\text{wrist}},d)_{t}\right)$ includes a language prompt $l$, a ego-centric image $I_{\text{ego}}$ (shared input), and an arm-specific wrist image $I_{\text{wrist}}$ with proprioception $d$ (arm-specific input). We train $\pi_{\text{single}}(A_{t}\mid o^{\text{single}}_{t})$ to predict the action chunk for one arm. For bimanual manipulation, the observation aggregates both right ($R$) and left ($L$) arm-specific input, $o^{\text{twin}}_{t}=\left((l,I_{\text{ego}})_{t},(I_{\text{wrist}}^{R},d^{R})_{t},(I_{\text{wrist}}^{L},d^{L})_{t}\right)$, and the policy $\pi_{\text{twin}}(A^{R}_{t},A^{L}_{t}\mid o^{\text{twin}}_{t})$ outputs a joint action chunk for right and left arms.
3.2 Training VLAs with Conditional Flow Matching
We aim for the VLA model to predict continuous robot actions from observations. Each observation $o_{t}$ is tokenized: language via a language tokenizer, images via a vision encoder, and proprioception via an MLP encoder. We append a learnable readout token $r_{t}$ to the observation token sequence, which signals the model to produce actions. These tokens are fed into the VLM backbone, and the embedding $h_{t}$ corresponding to the readout token $r_{t}$ is obtained from the final hidden state.
To enable continuous action prediction from VLM outputs, we attach an action head using conditional flow matching for both $\pi_{\text{single}}$ and $\pi_{\text{twin}}$. The action head $v_{\theta}(A_{t}^{\tau},h_{t},d_{t})$ is trained with the following flow-matching loss function:
$\mathcal{L}^{T}(\theta)=\mathbb{E}_{p(A_{t}\mid o_{t}),q(A_{t}^{\tau}\mid A_{t})}\lVert v_{\theta}(A_{t}^{\tau},h_{t},d_{t})-\mathbf{u}(A_{t}^{\tau}\mid A_{t})\rVert^{2},$ (1)
where $h_{t}$ is the VLM output hidden state, $d_{t}$ is proprioception. The objective trains the action head to predict the reference flow from a noised action chunk $A_{t}^{\tau}$ to the target action chunk $A_{t}$, where $\tau\in[0,1]$ denotes the flow matching timesteps. We adopt a simple linear Gaussian path $q(A_{t}^{\tau}\mid A_{t})=N(\tau A_{t},(1-\tau)I)$, which has demonstrated robust performance across domains.
Specifically, we first sample a noisy action $A_{t}^{\tau}=\tau A_{t}+(1-\tau)\epsilon$, where the noise $\epsilon\sim N(0,I)$ and the timestep $p(\tau)=Beta(\frac{0.999-\tau}{0.999};1.5,1)$, following $\pi_{0}$ (black2024pi_0). The action head $v_{\theta}(A_{t}^{\tau},h_{t},d_{t})$ and the VLM are jointly trained end-to-end to predict the reference flow $u(A_{t}^{\tau}\mid A_{t})=\epsilon-A_{t}$.
During inference, we sample actions using the forward Euler integration method. Starting from $A_{0}\sim N(0,I)$, we iteratively update the action using the learned flow $v_{\theta}(A_{t}^{\tau},h_{t},d_{t})$:
$A_{t}^{\tau+\delta}=A_{t}^{\tau}+\delta v_{\theta}(A_{t}^{\tau},h_{t},d_{t}),$ (2)
where we set the sampling step $n=10$ and use $\delta=\frac{1}{n}$. The entire model, including the vision encoder and VLM backbone, is trained in an end-to-end manner.
4 TwinVLA
TwinVLA is a modular architecture that transforms a pretrained SingleVLA into a coordinated bimanual policy through three core principles. First, we selectively duplicate modules from a pretrained SingleVLA to form a bimanual policy (Section˜4.1). We then introduce a joint attention mechanism to enable effective cross-arm coordination between the duplicated VLMs (Section˜4.2). Finally, we integrate a Mixture-of-Experts (MoE) layer for shared observations to enhance computational efficiency without sacrificing performance (Section˜4.3).
(a) Causal joint attention mask
(b) Transformer block of TwinVLA
Figure 3: (a) Causal attention mask for joint attention. It preserves causality while processing shared, left, and right inputs in parallel. (b) TwinVLA joint attention mechanism. The two VLMs share information, and the shared modality $(l,I_{\text{ego}})_{t}$ is further processed by MoE to more efficiently leverage both VLMs.
4.1 Single-Arm Policy Duplication
To construct TwinVLA from SingleVLA, we initialize the twin policies for the left and right arms by copying the pretrained SingleVLA. However, instead of duplicating the full model, we share the vision encoder and DiT (Peebles2022ScalableDM) action head while fully replicating the VLM. Each arm has its own lightweight proprioception encoder. This design yields a compact 1.31.3B-parameter model, comparable to the 1.21.2B-parameter RDT-1B, without significantly increasing computational cost.
Visual inputs are processed by the shared encoder, and each VLM produces readout tokens that are jointly decoded by the shared DiT. This design is motivated by the principle that general visual understanding (image encoding) and low-level motor control (action decoding) are largely embodiment-agnostic skills that can be effectively shared for both arms. In contrast, the VLM, which decides output action given encoded observation, is fully replicated to allow for specialized control.
4.2 Joint Attention for Cross-arm Fusion
We integrate arm-specific inputs using a joint attention mechanism inspired by Mixture of Transformers (MoT) (liang2024mixtureoftransformerssparsescalablearchitecture). As illustrated in Figure˜3(b), this is achieved by sharing only the self-attention layers across the VLM backbones, while other components like the projection networks operate independently for each arm. Unlike $\pi_{0}$ (black2024pi_0), which links VLM with an action head, we connect two VLMs directly. The detailed pseudocode of joint attention is presented in Algorithm˜1.
Causal joint attention mask. Effective joint attention requires appropriate attention masking. Standard LLMs use a lower-triangular attention mask for causal prediction. To support joint attention among the shared and arm-specific inputs, we designed the attention mask for TwinVLA as shown in Figure˜3(a). Specifically, we embed lower-triangular masks within each arm’s region while treating the shared modality as fully accessible. Each arm also attends to half of the other’s tokens, enabling symmetric cross-arm interaction without violating autoregressive constraints.
4.3 Mixture-of-Experts Integration
In TwinVLA, each VLM receives both shared and arm-specific inputs to produce single-arm actions. Feeding the shared modality $(l,I_{\text{ego}})_{t}$ redundantly to both VLMs is inefficient, as it increases token length and VRAM usage in a way that hinders the practical training of large models. To address this, we apply a Mixture-of-Experts (MoE) mechanism that routes the shared modality tokens between both VLMs, reducing token length while preserving representational power.
While the MoE effectively shares the feedforward networks (FFNs) between the twin VLM backbones, other key components like projection layers remain duplicated. To address this, we implement an effective form of task arithmetic (10.5555/3692070.3694018). We process an input through the corresponding component in each VLM and then average their outputs. This output-averaging technique functionally simulates a single shared layer without averaging the parameter itself. This is illustrated in the center of Figure˜3(b). Thanks to this change, we enable training the model with a batch size of $8$ on a single 4040GB GPU. Detailed implementation is provided in Appendix C.3.
Attention re-weighting. A potential side effect of introducing new arm-specific tokens is that the model’s learned attention patterns can be disrupted, shifting focus away from the pretrained shared modalities. To mitigate this and preserve the valuable pretrained knowledge, we re-scale the attention scores for the shared modality. This maintains pretrained modality importance, allowing the model to bypass an initial adaptation phase and focus directly on the target task—a benefit evidenced by a lower initial loss and converged loss during fine-tuning.
(a) Real-world tasks
(b) Simulation tasks
Figure 4: Experimental setups. (a) We evaluate TwinVLA on three real-world bimanual tasks using an Anubis robot. (b) We further analyze TwinVLA on a large suite of simulation tasks: 55 tasks in Tabletop-Sim and $50$ tasks in RoboTwin 2.0.
5 Experiments
In this paper, we propose TwinVLA to achieve strong bimanual manipulation performance with minimal bimanual data by fully leveraging a single-arm VLA pretrained on abundant single-arm data. Our empirical studies aim to answer the following questions:
  • How does TwinVLA compare to state-of-the-art methods across diverse bimanual tasks, without any bimanual pretraining (Sections˜5.2 and 5.3)?
  • How quickly can TwinVLA adapt to new bimanual tasks (Section˜5.4)?
  • Does TwinVLA retain core VLA properties—language-following and robustness to unseen scenes and instructions (Sections˜5.5 and 5.6)?
  • How much does each key design choice contribute to overall performance (Section˜5.7)?
5.1 Compared Methods
We evaluate TwinVLA against three bimanual manipulation policies, each representing a different point in the design space.
  • RDT-1B (liu2024rdt): This serves as our direct baseline. With a comparable size(1.21.2B vs. TwinVLA’s 1.31.3B parameters), it represents the standard monolithic approach that requires substantially larger resources (1.41.4M trajectories, $\sim$11,$440$ H100 days vs. 0.50.5M single-arm data, $\sim$$25$ H100 days).
  • $\bm{\pi_{0}}$ (black2024pi_0): We include this as a skyline, as this is 3.33.3B-parameter VLA trained on over 1010K hours of proprietary robot data. Our goal is to assess how closely TwinVLA can approach this performance ceiling with far greater efficiency.
  • Diffusion Policy (DP) (chi2023diffusion): This is a strong baseline method in low-data regime with 271271M parameters, used to demonstrate the crucial benefits of pretraining.
Figure 5: Success rates on real-world tasks. TwinVLA outperforms RDT-1B and DP on average. Moreover, TwinVLA shows comparable performance with $\pi_{0}$ while trained only on target data.
5.2 Real-World Experiments
Environment. For real-world experiments, we use a dual-arm robot, Anubis (anubis2025), as shown in Figure˜4(a). Anubis has two $6$ DoF arms with parallel-jaw grippers. The robot is equipped with two wrist-mounted cameras and a single ego-centric view camera.
Tasks. We design three long-horizon tabletop manipulation tasks which requires careful coordination and accurate motions: carrot to bag, brush to dustpan, and take towel off. We collect $50$ episodes for each task using absolute EEF control. We fine-tune all methods for each task, and evaluate them with $20$ rollouts per task.
Results. As presented in Figure˜5, our model, TwinVLA, significantly outperforms its direct competitor, RDT-1B. This result is particularly noteworthy given that TwinVLA was pretrained on a substantially smaller single-arm dataset (0.5M vs. RDT-1B’s 1.4M) with far less computational cost, demonstrating the data efficiency of our approach. The comparatively low performance of DP further underscores the critical importance of pretraining. Among all methods, $\pi_{0}$ ultimately demonstrated the best overall performance.
5.3 Simulation Experiments
RoboTwin 2.0. We use the RoboTwin 2.0 benchmark (chen2025robotwin), consisting of $50$ bimanual tasks. Adhering to the official evaluation protocol, we fine-tune a model per task with $50$ generated demonstrations and perform $100$ test rollouts under both “Easy” and “Hard” settings. For Easy tasks, test scenes match the training data, but the instructions are novel. The Hard tasks introduce variations in texture, object position, and height. For compared methods, we use the results reported from RoboTwin 2.0 (chen2025robotwin).
Tabletop-Sim. To assess dexterous scenarios beyond tasks in RoboTwin, we develop Tabletop-Sim [^1] Our simulation setup is similar to the concurrent work Aloha-Sim, released by Google DeepMind (aloha_sim_github)., a tabletop simulation environment based on dm_control (tunyasuvunakool2020) and assets from ALOHA2 (aloha2team2024aloha2enhancedlowcost) and GSO object dataset (DBLP:journals/corr/abs-2204-11918). We design $5$ representative tasks that require precise bimanual coordination. Specifically, we define four single-tasks and one multi-task: dish-drainer, handover-box, shoes-table, lift-box, and put X box into Y pot. In the “Hard” tasks, we vary background textures and objects. We collect $50$ episodes on each task using absolute EEF control, and fine-tune a model per task, and perform $500$ evaluation rollouts for both “Easy” and “Hard” settings.
Results. The results in Figure˜6 show the average success rates of TwinVLA and compared methods. DP, trained from scratch, shows the worst performance, highlighting the importance of pretraining. Once again, we observe that TwinVLA outperforms RDT-1B in most scenarios, except for the RoboTwin Hard tasks, and achieves comparable performance with $\pi_{0}$ by effectively leveraging single-arm data and modularity of bimanual manipulation. Notably, in Tabletop-Sim Easy tasks, TwinVLA even outperforms $\pi_{0}$, which is trained on an extensive corpus of high-quality bimanual pretraining data. This demonstrates TwinVLA’s advantages in scenarios demanding higher dexterity and significant bimanual coordination.
Figure 6: Average success rates for diverse bimanual tasks. Despite being pretrained solely on single-arm datasets, TwinVLA outperforms other methods except $\pi_{0}$.
5.4 Data Efficiency
Figure 7: Average success rates on the Tabletop-Sim Easy tasks. Models are evaluated after fine-tuning with 2020, $35$, and $50$ demonstrations.
TwinVLA exhibits data efficiency in two key aspects: pretraining and fine-tuning. For pretraining, it is efficient because it does not require supplemental bimanual data. For fine-tuning, it learns new tasks rapidly because its structural inductive bias facilitates the efficient transfer and application of its pretrained single-arm knowledge. We validate this efficiency in Tabletop-Sim Easy environment, comparing model’s average success rates with varying amounts of demonstration data. As illustrated in Figure˜7, TwinVLA exhibits a steep learning curve. Despite a modest start with $20$ demonstrations, it quickly surpasses the performance of RDT with just $50$ demonstrations, highlighting its exceptional data efficiency.
5.5 Language Following Evaluations
A known challenge is that fine-tuning VLMs on robotic data can degrade their ability to faithfully follow nuanced instructions. We therefore evaluate how effectively our model preserves this core capability in a multi-task setting. The “Put X box into Y pot” task involves $3$ box colors (X) and $2$ pot colors (Y), resulting in a total of $6$ possible instruction combinations. As observed in Figure˜8(a), TwinVLA outperforms both RDT-1B and $\pi_{0}$. We believe this performance stems from effectively preserving the knowledge acquired during single-arm pretraining through careful fine-tuning.
5.6 Policy Robustness
One of the advantages of VLAs is their robustness to unseen situations and novel language instructions, thanks to pretraining. As shown in Figure˜6, TwinVLA outperforms RDT-1B by $3.3\%$ even in the Hard setup of Tabletop-Sim, which involves different textures and objects.
The RoboTwin benchmark, both in the Easy and Hard setups, uses evaluation language instructions that are unseen during training. Here, TwinVLA again shows $7.48\%$ better performance than RDT-1B in the Easy setup. Although TwinVLA’s performance on the RoboTwin Hard tasks is $3.72\%$ lower than that of RDT-1B, it still outperforms a non-pretrained Diffusion policy by $9.38\%$. This result demonstrates that TwinVLA possesses sufficient robustness as a bimanual VLA, even without being pretrained on large-scale bimanual manipulation data.
(a) Language task results
(b) Ablation results
Figure 8: Language following task and ablation results. (a) We evaluate average success rates on the language following tasks in Tabletop-Sim. (b) Ablation studies in the real world and Tabletop-Sim Easy tasks.
5.7 Ablations
In this section, we conduct a sequential ablation study to analyze the cumulative impact of our key design choices on performance. Starting from the full TwinVLA model, we progressively remove each component in a specific order: first Attention Re-weighting, followed by MoE integration, and finally Joint Attention. This method reveals how performance degrades as each component of our architecture is stripped away. The results on our real-world and Tabletop-Sim Easy tasks are reported in Figure˜8(b).
Attention re-weighting. Removing the attention re-weighting mechanism (w/o Re-weighting) increased the initial fine-tuning loss by 40% and decreased final performance by 1.1% and 1.2% in simulation and real world, respectively. This demonstrates that our re-weighting strategy successfully mitigates the input distribution shift between pretraining and fine-tuning.
MoE integration. Building on the previous ablation, we next remove the MoE integration (w/o MoE). This additional change increased the token sequence length by 28% and increased VRAM usage by 21%, making VLA training more burdensome. Surprisingly, it also further decreases the success rate by 1.1% in simulation, suggesting that MoE integration eliminates redundant processing of shared inputs while maintaining the performance.
Joint attention. Lastly, removing the joint attention mechanism (w/o Joint attn) cause the most significant additional performance drops by 4.0% and 8.3% in simulation and real world, respectively. This impact is greater than other components, confirms that the joint attention is the critical mechanism for the bimanual manipulation requires coordination between both arms.
Effect of single-arm pretraining. As a separate, foundational experiment, we assess the role of pretraining by training a model from scratch without OXE dataset (Scratch). This resulted in a 4.6% performance drop in simulation and a stark 32.9% in real world. This result confirms that effective cross-arm coordination is essential for bimanual manipulation and validates joint attention as the critical mechanism for achieving it in our model.
Twin structure. While we have confirmed that joint attention effectively connects the two modules, a crucial question remains: how does this approach compare to a monolithic model that is inherently unified from the start? To answer this, we revisit our comparison against RDT-1B, a monolithic model of a comparable 1.2B parameter size. The results are telling: TwinVLA outperforms RDT-1B by 16.2% in the real world, 5.0% in simulation, and 25.1% in language-following tasks. This provides strong evidence that the inductive bias from the Twin Structure itself is highly beneficial for bimanual manipulation, validating our design choice over a monolithic approach.
6 Limitations
Since our work adapts a pre-trained single-arm VLA’s representations for bimanual tasks, a current limitation is that the model forgets its single-arm manipulation skills after fine-tuning. Future research into mechanisms that prevent this forgetting could address data scarcity by integrating diverse data, while also improving model explainability and better generalization capability to unseen tasks.
Moreover, the choice of action space is critical for VLA models, particularly for knowledge transfer. We adopt absolute end-effector (EEF) pose control because it provides an embodiment-agnostic representation essential for our single-arm transfer strategy. In contrast, joint positions are specific to a robot’s degrees of freedom (DOF), making them unsuitable for direct policy transfer. We believe that exploring relative absolute actions (Chi-RSS-24) or developing a shared representation across diverse embodiments could enable more efficient transfer.
7 Conclusion
In this paper, we introduce TwinVLA, a data-efficient VLA model for bimanual manipulation. TwinVLA provides a new perspective on solving bimanual manipulation under scarce bimanual data by leveraging abundant single-arm datasets. From a small amount of bimanual demonstration data, TwinVLA learns to coordinate two copies of a SingleVLA pretrained on large-scale single-arm data via our proposed joint attention. Through exhaustive experiments both in the real world and simulation, TwinVLA demonstrates its data-efficient learning of bimanual tasks compared to prior monolithic approaches. We believe this principle of bridging data availability gaps via leveraging modularity opens up exciting possibilities for other complex robotic domains, such as mobile manipulation, thereby broadening the impact of large-scale robotic learning.
Acknowledgments
This project was supported in part by Microsoft Research Asia and the Microsoft Accelerate Foundation Models Research (AFMR) grant program. This research was also supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korean Government (MSIT) (RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University); RS-2024-00436680, Global Research Support Program in the Digital Field Program), the Alchemist Project (RS-2024-00432143) funded by the Ministry of Trade, Industry & Energy (MOTIE, Korea), and the Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government (24ZR1100, A Study of Hyper-Connected Thinking Internet Technology by autonomous connecting, controlling, and evolving ways). LG Electronics provided the Anubis robot, which was used for the experiments.
Appendix
Appendix A SingleVLA: Efficient Single-Arm Policy Design and Pretraining
Figure 9: Overview of SingleVLA architecture design and pretraining method.
[…]
To ensure effective transfer to bimanual manipulation, it is crucial to choose an appropriate action space. Heterogeneous joint configurations across robots induce incompatible action spaces and complicate joint training. Prior work mitigates this with robot-specific decoders or high-dimensional zero-padded spaces (nvidia2025gr00tn1openfoundation; doshi2024scalingcrossembodiedlearningpolicy; octo_2023; black2024pi_0; liu2024rdt). Instead, we convert all actions into absolute end-effector (EEF) poses, providing a consistent, semantically meaningful representation across robots that naturally extends to bimanual control. For rotation, we adopt a 6D representation (8953486), which is well suited for neural network learning.
A.1 Pretraining
Table 1: SingleVLA pretraining datasets and sampling percentages.
Dataset Sample Percentage
RT-1 (rt12022arxiv) $24.49\%$
Kuka (filtered) (Yadav_Nagar_Shah_2024) $12.40\%$
BridgeV2 (walke2023bridgedata) $13.74\%$
Taco Play (rosete2022tacorl) $3.10\%$
Jaco Play (dass2023jacoplay) $0.50\%$
Viola (zhu2022viola) $1.00\%$
Berkeley Autolab UR5 (BerkeleyUR5Website) $1.28\%$
Stanford Hydra (belkhale2023hydra) $4.73\%$
Austin Buds (9695333) $0.22\%$
NYU Franka Play (cui2022play) $0.88\%$
FurnitureBench (heo2023furniturebench) $2.40\%$
Austin Sailor (nasiriany2022sailor) $2.33\%$
Austin Sirius (liu2022robot) $1.84\%$
DLR EDAN (shared control) (vogel_edan_2020; quere_shared_2020) $0.05\%$
UT Austin Mutex (shah2023mutex) $2.38\%$
Berkeley FANUC manipulation (zhu2023fanuc) $0.82\%$
CMU Stretch (bahl2023affordances; mendonca2023structured) $0.16\%$
BC-Z (filtered) (jang2021bc) $7.90\%$
FMB (doi:10.1177/02783649241276017) $7.40\%$
Dobb-E (shafiullah2023bringing) $1.50\%$
DROID (khazatsky2024droid) $10.70\%$
SingleVLA is pretrained on an OXE subset (0.5M trajectories); dataset composition and sampling rates appear in Table 1. We adopt the dataset loader from the OpenVLA (kim2024openvla) codebase and apply sampling according to the designated weights. Because some datasets (e.g., Kuka and BC-Z) include failed trajectories, we pre-process to retain only successful ones. All actions are converted to absolute EEF control with 6D rotations, resulting in a 10-Dimensional action space. We further apply frequency matching as described below.
Frequency matching.
Robotic datasets differ in control frequency, making fixed-length action-chunk prediction misaligned in real time. For example, a 20-step chunk spans $\sim 7$ seconds in RT-1 (rt12022arxiv) (3 Hz) but only $\sim 1.3$ seconds in DROID (khazatsky2024droid) (15 Hz). Mixing low-frequency data like OXE (oxe2024) with high-frequency datasets can degrade pretraining quality. Inspired by $\pi_{0}$-FAST (pertsch2025fastefficientactiontokenization), which uses DCT (1672377) to map $1$-second actions into a consistent space, we perform frequency matching via interpolation: all datasets are resampled to 20 Hz, improving temporal alignment and transfer to high-frequency bimanual tasks.
A.2 Hyperparameters and Compute
Table 2: Key hyperparameters for TwinVLA training.
Hyperparameter SingleVLA TwinVLA
Global batch size $256$ 88
Precision FP32/BF16 (mixed) FP32/BF16 (mixed)
Gradient clipping ($L_{2}$) $1.0$ $1.0$
Learning rate $\times 10^{-4} 1×1$$\times 10^{-4}$
LR scheduler cosine cosine
Warm-up ratio $0.01$ $0.05$
Total steps $120\text{k}$ $100\text{k}$
Optimizer AdamW AdamW
Weight decay $\times 10^{-5} 1×1$$\times 10^{-5}$
Adam $\epsilon$ $\times 10^{-8} 1×1$$\times 10^{-8}$
Vision backbone frozen true true
Image augmentation true false
[…]
A.3 SingleVLA VLM Ablation
We validate SingleVLA’s VLM choice in the LIBERO (liu2023libero) environment using several VLMs. The LIBERO actions are converted to absolute EEF 6D control. Due to computational limits, we directly fine-tune the pretrained VLM checkpoints on LIBERO (i.e., without additional pretraining on LIBERO). Each model is evaluated with $500$ rollouts per task suite under identical random seeds. Results are shown in Table 3.
Table 3: Performance of different VLMs on LIBERO.
VLM Spatial Object Goal Long Average
Qwen2VL-2B (wang2024qwen2vlenhancingvisionlanguagemodels) $0.4\% 88$$8.6\%$ $83.8\%$ $43.0\%$ $73.9\%$
InternVL2.5-1B (chen2025expandingperformanceboundariesopensource) $64.6\%$ $84.8\%$ $78.4\%$ $46.2\%$ $68.5\%$
Eagle2-1B (li2025eagle2buildingposttraining) $73.4\%$ $85.4\%$ $90.8\%$ $46.6\%$ $\mathbf{74.0\%}$
Although Qwen2VL is widely regarded as robust, Eagle2-1B achieves comparable or slightly better results while using roughly half the parameters and providing significantly faster inference. We therefore select Eagle2-1B as the VLM backbone for SingleVLA.
Table 4: Performance of pretrained SingleVLA on LIBERO.
Method Spatial Object Goal Long Average
SingleVLA (Eagle2-1B, no pretraining) $73.4\%$ $85.4\%$ $90.8\%$ $46.6\%$ $74.0\%$
SingleVLA (pretrained) $\mathbf{92.4\%}$ $\mathbf{94.5\%}$ $\mathbf{93.5\%}$ $\mathbf{63.7\%}$ $\mathbf{86.0\%}$
OpenVLA (kim2024openvla) $4.7\% 88$$8.4\%$ $79.2\%$ $53.7\%$ $76.5\%$
Octo (octo_2023) $78.9\%$ $85.7\%$ $84.6\%$ $51.1\%$ $75.1\%$
After pretraining SingleVLA with Eagle2-1B, we fine-tune it on LIBERO to assess single-arm capability. As shown in Table 4, the pretrained SingleVLA substantially improves performance and even surpasses the 7B model OpenVLA, indicating that the learned single-arm policy is both effective and sufficiently strong to benefit the bimanual policy.
Appendix B Implementation Details
Table 5: Training hyperparameters for baseline models.
Method # of params Learning rate Lr scheduler Batch size Training steps
TwinVLA 1.3B 1e-4 cosine 8 100k
RDT-1B 1.2B 1e-4 constant 8 100k
DP 271M 2e-5 cosine 8 100k
$\pi_{0}$ 3.3B 2.5e-5 cosine 8 100k
We use the official implementation of RDT-1B. Diffusion Policy and $\pi_{0}$ are evaluated via the public LeRobot release (cadene2024lerobot), with two modifications for a fair comparison. First, the LeRobot evaluation script normalized images differently from training; we corrected this to match the training pipeline.
All models are fine-tuned with the same number of steps and batch size so that the total number of training samples is consistent across methods. For learning rates, we began with each model’s default and tuned within a similar compute budget. In practice, defaults worked well for DP and RDT-1B. For $\pi_{0}$, we observed better final returns by slowing the cosine decay; we therefore extended the LR schedule from 30k to 100k steps.
Appendix C TwinVLA Details
C.1 Joint Attention
The joint attention in TwinVLA is fundamentally almost identical to the implementation in the Mixture-of-Transformers (MoT) (liang2024mixtureoftransformerssparsescalablearchitecture) study. While MoT has transformers for text, image, and speech inputs, in TwinVLA, the inputs for the left and right arms correspond to these.
Furthermore, MoT requires an operation to group mixed inputs by modality and then restore their original order. However, this process is unnecessary in TwinVLA because the inputs are fed in a fixed sequence: left arm, then right arm. The detailed computation process is shown in Algorithm˜1.
Algorithm 1 Joint Attention Computation
1:Let $x=(x_{1},\dots,x_{n})$ be the input sequence with $x_{i}\in\mathbb{R}^{d}$; let $m_{i}\in\{\text{left arm},\text{right arm}\}$ denote the modality of $x_{i}$.
2:Let $\mathcal{M}=\{\text{left arm},\text{right arm}\}$.
3:for each modality $m\in\mathcal{M}$ do
4:  $I_{m}\leftarrow\{i:m_{i}=m\}$ $\triangleright$ Indices for modality $m$
5:  $X_{m}\leftarrow\{x_{i}:i\in I_{m}\}$ $\triangleright$ Group tokens by modality
6:  $Q_{m}\leftarrow W_{Q}^{m}X_{m},\;K_{m}\leftarrow W_{K}^{m}X_{m},\;V_{m}\leftarrow W_{V}^{m}X_{m}$ $\triangleright$ Modality-specific projections
7:end for
[…]
10:for each modality $m\in\mathcal{M}$ do
11:  $O_{m}\leftarrow W_{O}^{m}A|_{I_{m}}$ $\triangleright$ Modality-specific output projection
12:  $H_{m}\leftarrow X_{m}+\mathrm{LayerNorm}^{m}_{\text{attn}}(O_{m})$ $\triangleright$ Residual & norm
13:  $F_{m}\leftarrow\mathrm{FFN}_{m}(H_{m})$
14:  $Y_{m}\leftarrow H_{m}+\mathrm{LayerNorm}^{m}_{\text{ffn}}(F_{m})$
15:end for
16:return $\{Y_{m}:m\in\mathcal{M}\}$
C.2 Attention Re-weighting
Figure 10: Due to the increased token length and softmax normalization, each VLM of TwinVLA refers to arm-specific inputs more than during pretraining, requiring the model to adapt.
Attention re-weighting is a technique we employ to improve the efficiency of adapting a pretrained SingleVLA into a bimanual TwinVLA. Constructing TwinVLA involves adding a second set of arm-specific modality tokens. During operation, input tokens are processed by their corresponding arm’s VLM backbone, pass through a joint attention layer, and then flow back to the individual VLMs. However, the softmax normalization within this joint attention layer presents a challenge. Although the total sequence length doubles, the number of tokens for shared inputs remains unchanged. Consequently, the proportion of attention allocated to these shared inputs is significantly diluted compared to the pretraining phase, creating a distribution shift for each VLM backbone’s inputs, as illustrated in Figure˜10.
Figure 11: By re-weighting the attention weights, we can make each VLM refer to each modality identically to its pretraining stage, resulting in no adaptation and a lower initial loss.
This discrepancy requires greater adaptation effort for TwinVLA during fine-tuning on bimanual tasks. To address this, we introduce a simple re-weighting trick immediately after the attention scores are calculated. Specifically, we double the attention weights corresponding to the shared modality tokens and then re-normalize all weights to sum to one. This adjustment effectively restores the proportional attention each VLM backbone assigns to the shared inputs, aligning it with the pretraining conditions (see Figure˜11). Applying this method reduced the initial fine-tuning loss by approximately 40%. While TwinVLA could learn bimanual manipulation without this technique, the required adaptation period would be substantially longer. This simple trick makes the process significantly more efficient and faster. We illustrate our implementation with simple pseudocode in Algorithm˜2.
Algorithm 2 Attention Re-weighting
Input: Attention weights $\mathbf{A}$, Scale factor $\alpha=2$
2:Output: Re-weighted attention weights $\mathbf{A}^{\prime}$
function ApplyReweighting($\mathbf{A}$, $\mathbf{c}$, $\alpha$)
4:  Create mask $\mathbf{M}$ $\triangleright$ Create a mask for arm-specific inputs
  $\mathbf{A_{reweighted}}\leftarrow\mathbf{A}\odot(\mathbf{M}+\alpha\cdot\neg\mathbf{M})$ $\triangleright$ Apply scaling to attention weights using the mask
6:  $\mathbf{A_{reweighted}}\leftarrow\text{Normalize}(\mathbf{A_{reweighted}})$ $\triangleright$ Normalize the new weights
  return $\mathbf{A}+(\mathbf{A_{reweighted}}-\mathbf{A})$ $\triangleright$ Return weights as a residual update for gradient flow
8:end function
C.3 MoE integration
To enable sharing of the shared inputs between the two-arm models, we duplicated the entire VLM transformer. This necessitates different strategies for sharing the FFNs and the other components. This section details the strategy used for each component of the transformer.
Feed-Forward Networks. To share FFNs, we adopt the common approach of using a gating-based MoE. In standard MoE, multiple FFNs are included within a transformer, and a gating mechanism activates a subset for each input. In TwinVLA, the two VLMs act as distinct FFN experts. Because shared inputs (e.g., egocentric views or language prompts) may have asymmetric relevance for each arm, the gating mechanism learns how much each FFN should contribute to processing the shared input. This approach is widely used and has been shown to improve training stability and preserve information more effectively than simple averaging (shazeer2017outrageously).
Other Components. Beyond FFNs, elements such as layer normalization and projection layers also require integration. For these, we apply task arithmetic (10.5555/3692070.3694018), merging the two VLMs via simple parameter averaging with weight $\lambda=0.5$. This extends MoE-style computation to the full transformer architecture.
Appendix D Real-World Robot Experiment Details
D.1 Initial Distribution
Figure 12: Initial distribution of each tasks in real-world.
To illustrate the diversity of initial configurations in our dataset, Figure˜12 shows an overlay of the first frames from all 50 demonstrations. For each demonstration, the position and orientation of the objects were randomized, resulting in a unique starting setup.
D.2 Quantitative Results
Table 6: Success rates for each model across all subtasks. The best overall performance is highlighted in bold. As $\pi_{0}$ is included as a skyline, as this is excluded from this direct comparison.
Task Subtask DP TwinVLA RDT-1B $\bm{\pi_{0}}$
Carrot to bag Pick up carrot $0.50$ $1.00$ $0.75$ $0.85$
Put carrot $0.20$ $0.70$ $0.40$ $0.65$
Close bag $0.15$ 0.65 $0.35$ 0.65
Brush to dustpan Move the brush $0.70$ $1.00$ $1.00$ $1.00$
Pick up the brush $0.65$ $1.00$ $1.00$ $1.00$
Put onto dustpan $0.35$ 0.80 0.40 0.80
Take towel off Dragging $0.40$ $0.90$ $0.80$ $0.95$
Half off $0.35$ $0.70$ $0.70$ $0.85$
Entirely off $0.20$ $0.55$ $0.60$ 0.65
We provide the quantitative results on real-world experiments in subtask-level detail in Table˜6. The results reveal the main bottleneck in each long-horizon task. The Carrot to bag task is challenging when inserting the carrot, which requires precisely opening the bag. The Brush to dustpan task’s bottleneck is the high-precision insertion of the brush into the dustpan. Lastly, in Take towel off rack, the final unfolding is difficult—unlike the simple initial steps—as it requires a successful switch between the arms. In the next subsection, we show qualitative results from these specific bottleneck phases.
D.3 Qualitative Results
Figure 13: Qualitative visualization of real world experiments.
Figure˜13 presents qualitative results highlighting challenging situations for each task. A check mark was used when the model succeeded with a probability above 0.5, an X mark for probabilities below 0.3, and an exclamation mark icon for intermediate cases.
  • Carrot to bag. $\pi_{0}$ showed the highest success rate, followed by TwinVLA, RDT, and DP. DP failed to interact meaningfully with the bag, especially struggling to grasp the cover properly. RDT failed to complete the task successfully, primarily due to its inability to accurately localize and grasp the bag’s opening.
  • Brush to dustpan. DP struggled either to grasp the brush itself or to successfully insert it. Interestingly, the RDT managed to grasp the brush well but lacked precision during the insertion. In this task, TwinVLA and $\pi_{0}$ demonstrated the same success rate.
  • Take towel off. DP mostly failed to pull the doll from a distant position toward the center, while the other models succeeded in pulling it to the center but showed differences in towel removal. Both RDT and $\pi_{0}$ tended to successfully remove one side of the towel and then easily remove the other side as well. In contrast, TwinVLA struggled with removing the remaining part and repeated the same action. This is likely because the longer action chunk length of RDT and $\pi_{0}$ helped them overcome the multimodality challenge.
D.4 Robot Hardware Spec
We conduct our real-world experiments using a custom-built robot named Anubis. The platform features a teleoperation system inspired by the Mobile ALOHA setup (fu2024mobile). Each arm has 6 DoF and is equipped with a parallel gripper and a wrist-mounted camera. At the center of the robot, an Intel RealSense camera is mounted on a height-adjustable mechanism, serving as the ego-centric view camera. Details are described in Table˜7. Anubis is equipped with a 3-wheel omni-directional base that supports planar locomotion; however, in this work, the mobility feature is not utilized.
Table 7: Anubis Robot Hardware Specifications.
Component Specification
Base Type 3-wheel omni-directional chassis
Mobility DOF 3 (X, Y, Yaw)
Arm DOF 2 × (6 DOF + gripper) = 14
Total Action Space 17 DOF
Wrist Cameras Intel RealSense D405
Gripper Parallel transparent gripper (hole design, ALOHA-style)
Power System 3 × Greenworks 40V 5.0Ah batteries (PC, wheels & leader/follower)
Frame 3D-printed custom components
Figure 14: The Anubis robot.
Appendix E Simulation Experiment Details
Figure 15: Task list of Tabletop-Sim.
E.1 Tabletop-Sim
To test bimanual policies in simulation, we developed Tabletop-Sim, a new benchmark specifically engineered to evaluate dexterous manipulation skills, in contrast to other benchmarks (mu2025robotwindualarmrobotbenchmark) that primarily focus on task diversity. The benchmark comprises four single-task environments and one multi-task setup. Our task selection was guided by the taxonomy in DexMimicGen (11127809), which categorizes bimanual tasks into: (1) parallel (two arms are doing separate tasks simultaneously), (2) coordinated (two arms are closely working together), and (3) sequential (one arm completes the task, and the other arm takes over) interactions. Using a custom controller similar to GELLO (wu2024gellogenerallowcostintuitive), we collected 50 demonstrations for each single-task and 60 for the multi-task environment.
The multi-task setup is a language-following task requiring the policy to place a specific box (out of three) into a designated pot (out of two) based on a language instruction. This task is designed to rigorously assess a model’s instruction-following capabilities, as Vision-Language-Action (VLA) models often disregard instructions after fine-tuning.
Furthermore, to evaluate policy robustness, we established two difficulty settings for the four single-tasks. The original tasks are designated as the Easy setting, while a Hard variant for each task incorporates challenging variations such as different textures, object models, and the presence of distractor objects. Figure˜15 presents snapshots of each task.
To ensure reproducibility and support future research, we will fully open-source this simulation and dataset.
E.2 Quantitative Results
This section describes the detailed results for the simulation tasks. The results for Tabletop-Sim are listed in  Table˜8, while the results for the RoboTwin 2.0 benchmark are in  Table˜9. For RoboTwin, the results for other baselines were referenced from the official benchmark results.
Although $\pi_{\textbf{0}}$ achieves the highest overall performance, this result is unsurprising considering its larger model size and pretraining dataset. Meanwhile, TwinVLA demonstrates consistently superior performance compared to RDT-1B, a model of a similar scale.
Table 8: Performance comparison on the Tabletop-Sim benchmark.
Tabletop-Sim
Dish drainer Handover box Lift box Shoes table Put X cube in to Y pot
Model Easy Hard Easy Hard Easy Hard Easy Hard
DP $0.686$ $0.590$ $0.180$ $0.086$ $0.100$ $0.006$ $0.028$ $0.260$ -
RDT-1B $0.810$ $0.780$ $0.694$ $0.508$ $0.300$ $0.076$ $0.660$ $0.192$ $0.555$
TwinVLA 0.954 0.836 $0.780$ 0.530 $0.452$ $0.044$ 0.848 $0.306$ 0.806
PI-0 $0.774$ $0.520$ 0.788 $0.444$ 0.512 0.136 $0.824$ 0.660 $0.792$
Table 9: Success rates of TwinVLA for $50$ bimanual tasks in RoboTwin 2.0.
Task Name Easy Hard Task Name Easy Hard
adjust bottle $0.97$ $0.35$ place can basket $0.40$ $0.00$
beat block hammer $0.77$ $0.10$ place cans plasticbox $0.47$ $0.08$
blocks ranking rgb $0.58$ $0.00$ place container plate $0.77$ $0.04$
blocks ranking size $0.03$ $0.00$ place dual shoes $0.18$ $0.03$
click alarmclock $0.33$ $0.01$ place empty cup $0.50$ $0.01$
click bell $0.58$ $0.13$ place fan $0.34$ $0.00$
dump bin bigbin $0.80$ $0.34$ place mouse pad $0.50$ $0.00$
grab roller $0.96$ $0.22$ place object basket $0.48$ $0.03$
handover block $0.17$ $0.00$ place object scale $0.06$ $0.00$
handover mic $0.84$ $0.02$ place object stand $0.20$ $0.02$
hanging mug $0.10$ $0.05$ place phone stand $0.34$ $0.02$
lift pot $0.87$ $0.07$ place shoe $0.48$ $0.04$
move can pot $0.45$ $0.05$ press stapler $0.62$ $0.26$
move pillbottle pad $0.32$ $0.02$ put bottles dustbin $0.08$ $0.04$
move playingcard away $0.61$ $0.35$ put object cabinet $0.39$ $0.16$
move stapler pad $0.11$ $0.00$ rotate qrcode $0.54$ $0.03$
open laptop $0.80$ $0.17$ scan object $0.11$ $0.04$
open microwave $0.03$ $0.01$ shake bottle horizontally $0.96$ $0.55$
pick diverse bottles $0.16$ $0.08$ shake bottle $0.93$ $0.58$
pick dual bottles $0.18$ $0.12$ stack blocks three $0.00$ $0.00$
place a2b left $0.27$ $0.05$ stack blocks two $0.26$ $0.00$
place a2b right $0.15$ $0.01$ stack bowls three $0.77$ $0.15$
place bread basket $0.11$ $0.03$ stack bowls two $0.84$ $0.11$
place bread skillet $0.20$ $0.01$ stamp seal $0.16$ $0.01$
place burger fries $0.67$ $0.13$ turn switch $0.25$ $0.15$
Average
Diffusion Policy $0.280$ $0.006$
RDT-1B $0.345$ $0.137$
TwinVLA $0.420$ $0.089$
$\pi_{0}$ $0.464$ $0.163$